We buy a bag of lime and cherry candies from a new manufacturer who uses candy wrappers colored red and green. The wrapper for each candy is selected probabilistically depending on the flavor. Notice that the corresponding probability model has three parameters  , e1, e2 as shown below. Suppose we unwrap N candies, of which c are cherries and l are limes. The wrapper counts are as follows: rc of the cherries have red wrappers and gc have green, while rl of the limes have red and gl have green. Find the maximum-likelihood values of the three parameters.

P(F=cherry) Flavor P(W=red | F) F |cherry 02 |lime Wrapper

, e1, e2
P(F=cherry) Flavor P(W=red | F) F |cherry 02 |lime Wrapper
Answer #1

0 = c/N

1=rc/(rc+ge) = r/(r +g)

02 rl/(rlgl) r/(rg)